Assume that the risk-free rate is 4.5% and the market risk premium is 5%. What is the required return for the overall stock market? Round your answer to one decimal place.

What is the required rate of return on a stock with a beta of 1.4? Round your answer to one decimal place.
Answer #1

market risk premium=market rate-risk free rate

market rate=4.5+5=9.5%

Required return=risk free rate+beta*(market rate-risk free rate)

=4.5+1.4*(9.5-4.5)

=11.5%